Suppose that the firm operates in a perfectly competitive market. the market price of its product is $2.the firm estimates its cost of production with the following cost production with the following cost function: C=100-6Q+2Q2

Based on this information, answer the following question.

A.what is the total fixed cost for 4 unit of output?

B.what is MC when the firm produces 4 unit of output?

C. What is the value of AFC at 4 units of output?

D.what is total variable cost for 4 unit of output?

E.What level of output should the firm produce to maximize its output?


1
Expert's answer
2022-03-23T18:32:32-0400

P=$2

"C=100-6Q+2Q^2"

TR=P×Q=2Q


A) "C=100-6Q-2Q^2"

"FC=TC-VC"

"FC=\$100"


B) "MC=\frac{dC}{dQ}=-6+4Q"

At Q=4

"MC=-6+4(4)=10"


C) "AFC=\frac{FC}{Q}"

At Q=4,

"AFC=\frac{100}{4}=25"


D) "TVC=TC-FC"

"TVC=100-6Q+2Q^2-100"

"TVC=-6Q+2Q^2"

At Q=4

"TVC=-6(4)+2(4)^2"

"TVC=8"


E) To maximize output, MR=MC

TR=2Q

MR=2

2=-6+4Q

"Q^*=2"